OpenStudy (anonymous):

2/3 x + 4= 3/5 x - 2 its a fraction

OpenStudy (anonymous):

\[\frac{2}{3}+4=\frac{3}{5}x-2\] do you know the LCD?

OpenStudy (anonymous):

Lcd?

OpenStudy (anonymous):

or GCD never mind ok we want to get rid of the fractions so what could we do to get rid of 2/3?

OpenStudy (anonymous):

Multiply 2/3 by 3/2 ?

OpenStudy (anonymous):

ok thats a start, but remember, we need to multiply something to the entire side is we multiplied by 3/2 to both sides, we would still have a fractions instead what would happen if you multiplied by 3 to the entire thing

OpenStudy (anonymous):

so it'd be the bottom number, 3, multiplying into 2/3 and cancel out the 3's?

OpenStudy (anonymous):

\[\frac{3}{2}(\frac{2}{3}x+4)=(\frac{3}{5}x−2)\frac{3}{2}\] actually if you multiplied 3/2 to both sides, it works out pretty well so you would be right

OpenStudy (anonymous):

cause 4 and 2 are both even :P

OpenStudy (anonymous):

oh ok, then what would i do after

OpenStudy (anonymous):

lets first look at what you would get once you multiplied by 3/2 \[x+6=\frac{9}{10}x-3\] we want to get rid of the fraction, any guesses?

OpenStudy (anonymous):

uhm.......? lol i dont understand

OpenStudy (anonymous):

you do know hte distributive property right? a(x+y)= ax+ay

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i just multiplied by 3/2 to get rid of the 2/3 fraction but whenever you do something to one side, you need to do it to the other you cant just multiply something by 3/2, you need to multiply the entire thing by 3/2

OpenStudy (anonymous):

ok i think i understand

OpenStudy (anonymous):

so now that we ahve x+6=9/10 x−3 what do you think we should multiply to the entire thing to get rid of the fraction?

OpenStudy (anonymous):

hint: try 10

OpenStudy (anonymous):

I dont understand again this is really confusing haha

OpenStudy (anonymous):

we want to get rid of the fraction 9/10 one way of doing so would be multiplying 10 to it 9/10*10=?

OpenStudy (anonymous):

\[\frac{9}{10}*10=???\]

OpenStudy (anonymous):

\[10(x+6)=(\frac{9}{10} x−3)10\] \[10x+60=9x−30\] solve for x